OpenStudy (phoenixfire):

Prove that if n-2 is divisible by 4 then n^2 - 4 is divisible by 16.

OpenStudy (phoenixfire):

For a given integer.

OpenStudy (anonymous):

Say that n is 26 so 26-2=24 and 24 is divisible by 4. 26x 2-4=52-4=48 and 48 is divisible by 16

OpenStudy (zzr0ck3r):

hehe I wish, sec

OpenStudy (zzr0ck3r):

I would do contradiction

OpenStudy (phoenixfire):

\[\forall{n}\in \mathbb{Z} : 4|n-2 \rightarrow 16|n^2-4\] I believe that's the correct notation.

OpenStudy (zzr0ck3r):

yeah

OpenStudy (zzr0ck3r):

do you need to show for all n?

OpenStudy (zzr0ck3r):

nm i c

OpenStudy (phoenixfire):

I need to show the proof.

OpenStudy (zzr0ck3r):

ok assume n-2= 4k for some k in Z then n = 4k+2 then n^2-4 = (4k+2)^2 - 4 = 16k^2 + 16k +4-4 = 16(k^2+k) since k^2+k is in Z 16|n^2-4

OpenStudy (zzr0ck3r):

sorry direct proof was fast I think

OpenStudy (phoenixfire):

Yeah, they wanted Direct Proof. so since n^2 - 4 = 16k the (k^2+k) in 16(k^2+k) doesn't matter, the rest match. that's what was confusing me.

OpenStudy (zzr0ck3r):

yeah 16 | (16* any integer)

OpenStudy (phoenixfire):

Thanks for the help.

OpenStudy (zzr0ck3r):

np